OpenStudy (anonymous):

write (27a^-9)^-2/3 in simplest form please help me ive been stuck on this for like 20 minutes now

OpenStudy (anonymous):

a^18/2187

OpenStudy (anonymous):

\[(27a^{-9})^{-\frac{2}{3}} = 27^{-\frac{2}{3}}(a^{-9})^{-\frac{2}{3}} = (3^3)^{-\frac{2}{3}}(a^{-9})^{-\frac{2}{3}}\] \[=3^{-3\frac{2}{3}}a^{-9\frac{-2}{3}} = 3^{-2}a^{6} = \frac{1}{3^2}a^6= \frac{1}{9}a^6\]

OpenStudy (anonymous):

@amistre64 Could you please validate my solution? Thank you in advance!

OpenStudy (anonymous):

I think my solution is valid if a is positive.

OpenStudy (amistre64):

even if a is negative its fine

OpenStudy (anonymous):

Thank you very much!